OpenStudy (anonymous):

Find the solutions to the system. y = x² – 5x + 2 y = –6x + 4  (1 point)

OpenStudy (anonymous):

@helpme1.2

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

PLEASE

OpenStudy (anonymous):

ill fan whoever helps

OpenStudy (anonymous):

So to solve this, you can substitute one equation into the other because they are both equal to y. So: x² – 5x + 2 = –6x + 4 Move everything to one side: 0= -6x + 4 - x^2 +5x -2 You can simplify this to: 0= -x^2 -x -2 Does this make sense so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):